Given two events G and H, the probabilities of each occurring are as follows: P(G) = 0.29; P(H) = 0.43; P(H AND G) =0.08. Using this information:
Find P(H OR G).
Type answer with 0 in front of decimal to 2 places. If answer i

Answer by ikleyn(52814)

.
P(H OR G) = P(H) + P(G) - P(H AND G) = 0.29 + 0.43 - 0.08.


Use your calculator.